OpenStudy (anonymous):

PLEASE HELP DUE TOMORROW What is the half life of each of these radioactive isotopes? A) a radioactive isotope decreased to one-half its original amount in 18 months. B) a radioactive isotope decreased to one-fourth its original amount 100 years.

OpenStudy (abhisar):

Hey there! What do you think it should be?

OpenStudy (abhisar):

Basically, half life is the time taken for the radioactivity of a specified isotope to fall to half its original value.

OpenStudy (anonymous):

honestly i am trying but i havent an inlking of the answer

OpenStudy (abhisar):

Read the definition i just told you, it says half life is the time taken to reduce to half of its original value. Now see what they say in Q#1, they say that the radioactive isotopes got reduced to half in 18 months. So what should be the half life of this isotope?

OpenStudy (anonymous):

6 months

OpenStudy (anonymous):

oh 18 months

OpenStudy (abhisar):

Yes, that's \(\huge \checkmark\)

OpenStudy (anonymous):

thx and number 2

OpenStudy (anonymous):

50 years

OpenStudy (anonymous):

am i right on 50 years and Thank you so much for the help can you help with this to please i need to determine wheter its a chemical or nuclear reaction occuring. A) when two compounds are combined heat is released. B) a sample of galllium-68 is reduced to one-half of its original amount in 68.3 minutes. C) Radium-226 decays to radon 222 D) A spark of energy is used to begin the combustion of methane gas. E) Hydrogen nuclei are fused in the sun to make helium atoms.

OpenStudy (abhisar):

recognize 1/4 as a fraction associated with 2 half-lives (from (1/2)2 = 1/4) \(\therefore\) Half life X 2=100 => Half life = 100/2=50

OpenStudy (abhisar):

Ask, one question per post. Repost your next question in a new thread \(\huge \ddot \smile\)

OpenStudy (abhisar):

ok, do that from next question....let's do this one here only.. Let's talk one by one. What do you think about A ?

OpenStudy (anonymous):

chemical

OpenStudy (abhisar):

Yes, that's correct. Whenever two \(\bf Compounds\) combine and heat is released or absorbed it is chemical reaction. If heat is released then it is called as exothermic chemical reaction and if energy is absorbed then it is called as endothermic chemical reaction... Next

OpenStudy (anonymous):

nuclear

OpenStudy (abhisar):

Yes, radioactivity is nuclear reaction. Next

OpenStudy (anonymous):

chemical

OpenStudy (abhisar):

Sorry, next is C. It's decay, it's also radioactivity and hence it is nuclear reaction and not chemical reaction.

OpenStudy (abhisar):

Got it?

OpenStudy (anonymous):

d is che,mical

OpenStudy (abhisar):

YEs, D is combustion as I told you, it's exothermic chemical reaction. Next

OpenStudy (anonymous):

e is nuclear

OpenStudy (abhisar):

Very Good! Last one is Fusion reaction which is a nuclear reaction. this reaction is the source of sun's energy.

OpenStudy (anonymous):

thx for all the help

OpenStudy (abhisar):

Ur most welcome :)